OpenStudy (anonymous):

After you graph a slanted asymptote, do you make a point on where it crosses in the x axis?

OpenStudy (amistre64):

depends on how you are sketching the line ....

OpenStudy (amistre64):

a sketch simply has at least a minimal amount of information to avoid ambiguity 2 defined reference points and the line between them are required

OpenStudy (amistre64):

in many textbook cases, the 2 points are easily defined as the y and x intercepts

OpenStudy (amistre64):

for example: take the line y=4x+ 3 when x=0, y=3 when x=-3/4, y=0 

OpenStudy (amistre64):

but any 2 points of reference work: y = 4x + 3 when x=0, y=3 when x=1, y=7 

OpenStudy (amistre64):

the tic on the x axis is not needed, but more information is not bad ... its less information that gets us in trouble